OpenStudy (anonymous):

4x^2 +21x=-5 solve for x

OpenStudy (anonymous):

4x^2+21x+5=0 (4x+1)(x+5)=0

sam (.sam.):

4x^(2)+21x=-5 4x^(2)+21x+5=0 (x+5)(x+(1)/(4))=0 (x+5)(4x+1)=0 x=-5,-1/4

OpenStudy (apoorvk):

simplify and bring all terms to one side then split the middle term and factorise

OpenStudy (anonymous):

x= -5, -1/4